# LLM-Enhanced Causal Discovery in Temporal Domain from Interventional Data

### v1

@@ Abstract.
In the field of Artificial Intelligence for Information Technology Operations (AIOps), causal discovery is pivotal for operation and maintenance of graph construction, facilitating downstream industrial tasks such as root cause analysis. Temporal causal discovery, as an emerging method, aims to identify temporal causal relationships between variables directly from observations by utilizing interventional data. However, existing methods mainly focus on synthetic datasets with heavy reliance on interventional targets and ignore the textual information hidden in real-world systems, failing to conduct causal discovery for real industrial scenarios. To tackle this problem, in this paper we propose to investigate temporal causal discovery in industrial scenarios, which faces two critical challenges: 1) how to discover causal relationships without the interventional targets that are costly to obtain in practice, and 2) how to discover causal relations via leveraging the textual information in systems which can be complex yet abundant in industrial contexts. To address these challenges, we propose the RealTCD framework, which is able to leverage domain knowledge to discover temporal causal relationships without interventional targets. Specifically, we first develop a score-based temporal causal discovery method capable of discovering causal relations for root cause analysis without relying on interventional targets through strategic masking and regularization. Furthermore, by employing Large Language Models (LLMs) to handle texts and integrate domain knowledge, we introduce LLM-guided meta-initialization to extract the meta-knowledge from textual information hidden in systems to boost the quality of discovery. We conduct extensive experiments on both simulation and real-world datasets to show the superiority of our proposed RealTCD framework over existing baselines in discovering temporal causal structures.
@@
journalyear: 2024submissionid: 397ccs: Computing methodologies Causal reasoning and diagnosticsccs: Computing methodologies Temporal reasoning
@@ 1. Introduction
Temporal causal discovery, as an emerging approach, aims to directly identify temporal causal relationships between variables based on observational data, with the utilization of interventional data. This group of methods have gained significant attention in recent years due to their promising potential to uncover causal dependencies in dynamic systems. For example, Brouillard et al. (2020) and Li et al. (2023) employ temporal causal discovery methods to leverage various types of interventional data and have achieved remarkable progress in discovering the underlying temporal causal relationships.
@@
  • How to discover causal relations via leveraging the textual information in systems which can be complex yet abundant in industrial contexts?
To address these challenges, we propose the RealTCD framework, which is able to leverage the textual information from real-world systems to discover temporal causal relationships without interventional targets. Specifically, we first develop a score-based temporal causal discovery method that learns the underlying causal relationship without interventional targets through strategic masking and regularization. We impose regularizations on both the adjacency matrix and interventional family within the context of the regularized maximum log-likelihood score, as well as and optimize them in a joint manner. In this way, the costly interventional targets are not required for boarder applications in real-world industrial scenarios. Subsequently, by leveraging Large Language Models (LLMs) to handle texts, we introduce LLM-guided meta-initialization that infers and initializes the inherent causal structures from the textual information in systems for the aforementioned discovery process, which incorporates the domain knowledge while upholding the theoretical integrity of temporal causal discovery. Extensive experiments on both simulation and real-world datasets demonstrate the superiority of our RealTCD framework over existing baselines in terms of discovering temporal causal structures. The deeper analyses also show that our method can effectively discover the underlying temporal causal relationship without interventional targets in industrial scenarios. In summary, our main contributions are as follows:
@@
  • We propose the RealTCD framework, including two specially designed modules, i.e., score-based temporal causal discovery and LLM-guided meta-initialization, which is able to leverage the textual information in systems to discover temporal causal relationships without interventional targets in industrial scenarios.
@@
2. Related Work
2.1. Causal Discovery in Temporal Domain
Existing theoretical evidence confirms the fact that when interventional data is readily available, it greatly amplifies the process of identifying underlying causal structures (Eberhardt, 2012; Tian and Pearl, 2013; Gao et al., 2022; Zhang et al., 2023e). However, the expansion of research in this particular domain has been rather restricted due to substantial obstacles rooted in designing intervention experiments as well as gathering the requisite data.  Jaber et al. (2020) took up the challenge of learning causal graphs entailing latent variables, derived from a combination of observational and interventional distributions where the interventional aims are yet to be known. To address this concern, the study proposed an approach utilizing a $\Psi$-Markov property.  Addanki et al. (2020) put forth a randomized algorithm, characterized by $p$-colliders, that recovers the comprehensive causal graph, whilst also minimizing the intervention expense. Further, an adaptable method for causality detection,  Brouillard et al. (2020), notably benefits from different categories of interventional data and incorporates refined neural architectures like the normalizing flows, operating under continuous constraints.  Li et al. (2023) attempt to discover the temporal cause from observation data, while the intervention labels, which are hard to obtain in real-world scenarios, are indispensable for the proposed algorithm, limiting its application in practice. In comparison, in this paper, we focus on the setting of discovering the temporal causal relationship without intervention labels.
2.2. LLM for Causal Discovery
Recently, there have been several works about utilizing large language models (LLMs) for causal discovery (Jiralerspong et al., 2024; Vashishtha et al., 2023a; Bhattacharjee et al., 2023; Gui and Toubia, 2023; Kasetty et al., 2024; Liu et al., 2024; Willig et al., 2022; Takayama et al., 2024; Ban et al., 2023a; Cohrs et al., 2023) and dynamic graph (Zhang et al., 2023d; Wang et al., 2017, 2019; Zhang et al., 2023a).  Ban et al. (2023b) explore the use of LLMs as advanced causal architects to discover causal relationships from data and highlight the potential of LLMs in transforming traditional query tools into powerful tools for causal inference.  Chan et al. (2023) focus on evaluating the performance of ChatGPT in capturing sentence-level temporal, causal, and discourse relations.  Chen et al. (2023) explore how incorporating LLMs can enhance causal discovery methods by reducing biases and improving the accuracy of learned causal structures and propose to mitigate prior errors in causal structure learning by leveraging LLM-driven prior knowledge.  Kıcıman et al. (2023) explore the role of LLMs in advancing causal discovery, counterfactual reasoning, and actual causality tasks, and highlight the potential of large language models in opening new frontiers for causal reasoning.  Long et al. (2023) investigate the question of whether large language models can build causal graphs, and explore whether GPT-3 in understanding the causal relationships between variables in the medical context. To the best of our knowledge, we are the first to utilize LLMs in the field of temporal causal discovery.
@@ 4. Method
Figure 1. The framework of our proposed method RealTCD. Given the system textual information and temporal data without interventional targets, the LLM-guided Meta-Initialization module leverages LLMs to extract the domain knowledge and obtain the potential causal relations as the initialization adjacency matrix $M_{0}^{\mathcal{G}}$. Then, the Score-based Temporal Causal Discovery module utilizes an augmented Lagrangian process to optimize the score for unknown interventional targets under constraints, where the $\Lambda_{0}$ is initialized with $M_{0}^{\mathcal{G}}$. In this way, the proposed RealTCD leverages the system textual information to discover temporal causal relationships without interventional targets.
@@ 4.1.1. Raw data
The raw data we used in our method as shown in Figure 1 refers to a set of input temporal data that includes both observational and interventional data. Observational data refers to standard data without any interventions or anomalies, while interventional data contains interventions as discussed in Section 3.2, or represents abnormal data in real-world datasets. The colored columns in interventional data represent the time series of intervened (or anomalous) variables. This data encompasses $d$ different measurement points in a system over $p$ time-lagged terms, as described in Section 3.1. Additionally, there may be multiple sets of interventional data, each corresponding to different interventional targets or anomalies.
@@ 4.1.3. Score for unknown interventional targets
After that, we assume that the interventional targets are unknown. To still be able to utilize the special information from interventional data, we propose to jointly optimize the adjacency matrix and interventional family as well as the NN’s parameters, thus, reaching the two optimization goals simultaneously. To be specific, a regularization term for the interventional family is added to the above score, and we form the score for unknown interventional targets:
@@
where $|\mathcal{I}|=\sum_{q=1}^{Q}|\mathcal{I}_{q}|$ counts the total number of intervened nodes in data.
@@ 4.1.4. Maximize the score
Moreover, as for the gradient of the score w.r.t. $\alpha_{ij}$ and $\beta_{qj}$, following the general dealing method in continuous optimization for causal discovery (Kalainathan et al., 2022; Ng et al., 2022), we estimate $\Lambda$ and $\Gamma$ by Straight-Through Gumbel estimator, which means that Bernoulli samples are used in the forward pass and Gumbel-Softmax samples are used in the backward pass.
Overall, the learnable parameters in the process are $\phi$,$\Lambda$, $\Gamma$, and the estimated adjacency matrix reflecting temporal causal relations is $\sigma(\Lambda)$ and the estimated potential interventional family is $\sigma(\Gamma)$
@@ 4.2. LLM-guided Meta Initialization
Through the list of tuples LLM offered to us, we then construct the adjacency matrix $M_{0}^{\mathcal{G}}$ to denote the causal structure learned from LLM. Note that, these results may not be the certainly correct causality conform to a causal definition, but a range of potential causality based on the domain information. We further design two modules to incorporate the meta-initialization information.
@@
2:  Obtain the causal results from LLM and transfer into matrix $M_{0}^{\mathcal{G}}$
@@
7:     Get the sub solution $\phi_{t}^{*},\Lambda_{t}^{*},\Gamma_{t}^{*}$, and initialize $\phi_{t+1},\Lambda_{t+1},\Gamma_{t+1}$ by them
@@
11:  Form the causal graph by adding an edge whenever $\sigma(\Lambda)>0.5$ is acyclic
Guide temporal causal discovery from data
Extract weight
@@ 5.1.1. Baselines
  • NeuralGC (Tank et al., 2021): Since NeuralGC only learns contemporaneous relationships, we use $W_{\mathrm{full}}$ defined as below to compress both contemporaneous and time-lagged causal relationships learnt from RealTCD into a single metric for comparison with NeuralGC, which does not differentiate between these 2 types.
@@
The distinction in method selection across different settings was intentional and aligned with the capabilities of each algorithm: Since PCMCI and DYNOTEARS are designed for causal structure learning from observational data and can not incorporate interventional data, we included them in the ”Unknown” interventional targets setting. As for TECDI, we use the same data with RealTCD , i.e. both observational and interventional data (or abnormal data in real datasets), to run it, since it can utilize the interventional data. However, for other baseline models that are unable to bring in interventional data, we have ensured a fair comparison to some extent by keeping the overall sample size consistent across all models, while using only observational data (or normal data in real datasets) for the baseline models.
@@ 5.1.2. Synthetic datasets
Before training, all data are normalized by subtracting the mean and dividing by the standard deviation. We experimented on two simulated datasets: Dataset 1 contains 5 nodes, their 1 time-lagged variables and 5 different interventional targets, each of which covers a single different node. Dataset 2 contains 10 nodes, their 1 time-lagged variables and 10 different interventional targets, each of which covers a single different node. We initially chose to limit our datasets to a time delay of 1 to facilitate the evaluation and presentation. However, it is indeed feasible to increase the time delay of the data.
5.1.3. Real-world data center datasets
@@
In contemporary data centers, the stability of IT equipment operation is paramount. To this end, sophisticated air conditioning systems are employed to manage the heat generated by the equipment, thereby sustaining a consistent indoor temperature. As illustrated in Figure 2, a standard data center room is methodically organized with rows of equipment (labeled A, B, C, D, and so forth). These rows are strategically separated by physical barriers, effectively isolating adjacent cabinet rows to impede the intermingling of hot and cold air streams. Computer Room Air Conditioners (CRACs), strategically positioned on both sides of the room, facilitate a closed-loop configuration, essential for preserving a stable environmental condition. The deployment of multiple sensors within the cold aisle is a critical aspect of this setup, providing real-time temperature data. This continuous monitoring is crucial for ensuring the uninterrupted delivery of cold air, thereby allowing for prompt adjustments that are essential for the consistent operation of IT equipment.
@@
The data used in this study was obtained from a specific data center at Alibaba. It covers monitoring data from a cooling system of a particular room from January 1st, 2023 to May 1st, 2023, and includes 38 variables in total. These variables comprise 18 cold aisle temperatures from sensors and 20 air conditioning supply temperatures from CRACs. We collected several time series of these 38 variables during normal as well as abnormal states. For the latter, data was sampled within 20 minutes of the occurrence of the abnormality, with each sampling interval being 10 seconds. Anomaly points were identified by learning the normal distribution range from historical data, using the $n$-$\sigma$ method. Any data points that fall outside of the $n$-$\sigma$ range (e.g., 3 to 5) of it selves are extracted as anomaly time points.
@@ 5.1.4. Evaluation metrics
We leverage the following two main metrics to evaluate the performance of the proposed method on learning causal graph: i) the structural Hamming distance (SHD), which calculates the number of different edges (either reversed, missing or redundant) between two DAGs; ii) the structural interventional distance (SID), which represents the difference between two DAGs according to their causal inference conditions (Peters and Bühlmann, 2015).
@@
Given the absence of ground truth DAGs in real case, we employ 4 types of performance metrics based on prior knowledge to assess the algorithms’ efficacy in learning causal graphs. These metrics mainly consider the physical location relationships within the room, shown in figure 2. Specifically: i) A2C True counts the correctly identified causal edges from air conditioning supply temperatures (A) to the temperatures of adjacent cold aisles (C), assuming that A influences C. ii) A2A True counts the correctly identified causal edges between adjacent air conditioning units, assuming mutual influences among them. iii) C2C True counts the correctly identified causal edges between temperatures in the same column of cold aisles, assuming direct interactions among them. iv) C2A False counts the incorrectly identified causal edges from cold aisle temperatures (C) to air conditioning supply temperatures (A), as such causality is implausible given that downstream variables (C) cannot influence upstream variables (A).
@@ 5.2.1. On synthetic datasets
The results on synthetic datasets are reported in Table 2. We can find that on both datasets, our method achieves significantly better results than baseline models on both metrics SHD and SID of the overall structures. The standard deviation of each metric is relatively small as well. Especially for the dataset with 10 nodes, our improvement over other baselines is more obvious, indicating that our method indeed has more advantages when dealing with a large number of variables, making the optimization more focused and effective. RealTCD consistently achieves better performance in each setting. Specifically, the ”Known” targets setting performs better than the ”Unknown” targets setting, which is reasonable since the former employs ground-truth interventional target labels for training. However, it is always impossible to obtain ground-truth interventional targets in real-world applications, making the proposed RealTCD for the setting of interventional data with unknown targets highly practical.
@@
Figure 3 and Figure 4 show example results of the temporal causal DAG and interventional targets simultaneously obtained through RealTCD. In Figure 4, each row represents a contemporary node, while each column in the matrix is a regime corresponding to an interventional family, where the first column represents the observational data, and the following 5 columns represent 5 different interventional data respectively. The pink cells are the learned/ ground truth interventional targets in each regime. It shows that our method is able to find the correct temporal causal graph and interventional targets at the same time.
Figure 3. A showcase of the DAG results on synthetic data.
Figure 4. A showcase of the interventional targets results on synthetic data.

### v2

# RealTCD: Temporal Causal Discovery from Interventional Data with Large Language Model

@@ Abstract.
In the field of Artificial Intelligence for Information Technology Operations, causal discovery is pivotal for operation and maintenance of systems, facilitating downstream industrial tasks such as root cause analysis. Temporal causal discovery, as an emerging method, aims to identify temporal causal relations between variables directly from observations by utilizing interventional data. However, existing methods mainly focus on synthetic datasets with heavy reliance on interventional targets and ignore the textual information hidden in real-world systems, failing to conduct causal discovery for real industrial scenarios. To tackle this problem, in this paper we investigate temporal causal discovery in industrial scenarios, which faces two critical challenges: how to discover causal relations without the interventional targets that are costly to obtain in practice, and how to discover causal relations via leveraging the textual information in systems which can be complex yet abundant in industrial contexts. To address these challenges, we propose the RealTCD framework, which is able to leverage domain knowledge to discover temporal causal relations without interventional targets. We first develop a score-based temporal causal discovery method capable of discovering causal relations without relying on interventional targets through strategic masking and regularization. Then, by employing Large Language Models (LLMs) to handle texts and integrate domain knowledge, we introduce LLM-guided meta-initialization to extract the meta-knowledge from textual information hidden in systems to boost the quality of discovery. We conduct extensive experiments on both simulation datasets and our real-world application scenario to show the superiority of our proposed RealTCD over existing baselines in temporal causal discovery.
@@
copyright: acmlicensedjournalyear: 2024submissionid: 849ccs: Computing methodologies Causal reasoning and diagnosticsccs: Computing methodologies Temporal reasoning
@@ 1. Introduction
Temporal causal discovery, as an emerging approach, aims to directly identify temporal causal relationships between variables based on observational data, with the utilization of interventional data. This group of methods has gained significant attention in recent years due to their promising potential to uncover causal dependencies in dynamic systems. Brouillard et al. (2020) and Li et al. (2023) employ temporal causal discovery methods to leverage various types of interventional data and have achieved remarkable progress in discovering the underlying temporal causal relationships.
@@
  • How to discover causal relationships via leveraging the textual information in systems which can be complex yet abundant in industrial contexts?
To address these challenges, we propose the RealTCD framework, which is able to leverage the textual information from real-world systems to discover temporal causal relationships without interventional targets. Specifically, we first develop a score-based temporal causal discovery method that learns the underlying causal relationship without interventional targets through strategic masking and regularization. We impose regularizations on both the adjacency matrix and interventional family within the context of the regularized maximum log-likelihood score, and optimize them in a joint manner. In this way, the costly interventional targets are not required for boarder applications in real-world industrial scenarios. Subsequently, by leveraging Large Language Models (LLMs) to handle texts, we introduce LLM-guided meta-initialization that infers and initializes the inherent causal structures from the textual information in systems for the aforementioned discovery process, which incorporates the domain knowledge while upholding the theoretical integrity of temporal causal discovery. Extensive experiments on both simulation and real-world datasets demonstrate the superiority of our RealTCD framework over existing baselines. Deeper analyses also show that our method can effectively discover the underlying temporal causal relationships without interventional targets in industrial scenarios. In summary, our main contributions are as follows:
@@
  • We propose the RealTCD framework, including two specially designed modules: score-based temporal causal discovery and LLM-guided meta-initialization, which is able to leverage the textual information in systems to discover temporal causal relationships without interventional targets in industrial scenarios.
@@ 3. Method
Figure 1. The framework of our proposed method RealTCD. Given the system textual information and temporal data without interventional targets, the LLM-guided Meta-Initialization module leverages LLMs to extract the domain knowledge and obtain the potential causal relationships as the initialization adjacency matrix $M_{0}^{\mathcal{G}}$. Then, the Score-based Temporal Causal Discovery module utilizes an augmented Lagrangian process to optimize the score for unknown interventional targets under constraints, where the $\Lambda_{0}$ is initialized with $M_{0}^{\mathcal{G}}$. In this way, the proposed RealTCD leverages the system textual information to discover temporal causal relationships without interventional targets.
@@ 3.1.1. Raw data
Raw data we used as shown in Figure 1 refers to a set of input temporal data including both observational and interventional data. Observational data refers to standard data without interventions or anomalies, while interventional data contains interventions discussed in Section 2.2, or represents abnormal data in real-world datasets. The colored columns represent time series of ground truth interventional targets that are unknown in prior. There may be multiple sets of interventional data, each corresponding to different interventional targets or anomalies.
@@ 3.1.3. Score for unknown interventional targets
Then, we assume the interventional targets are unknown. To still be able to utilize the special information from interventional data, we propose to jointly optimize the adjacency matrix and interventional family as well as the NN’s parameters, thus, reaching the two optimization goals simultaneously. To be specific, a regularization term for the interventional family is added to the above score, and we form score for unknown interventional targets:
@@
where $|\mathcal{I}|=\sum_{q=1}^{Q}|\mathcal{I}_{q}|$ counts the total number of intervened nodes.
@@ 3.1.4. Maximize the score
Moreover, as for gradient of the score w.r.t. $\alpha_{ij}$ and $\beta_{qj}$, following the general dealing method in continuous optimization for causal discovery (Kalainathan et al., 2022; Ng et al., 2022), we estimate $\Lambda$ and $\Gamma$ by Straight-Through Gumbel estimator, which means that Bernoulli samples are used in forward pass and Gumbel-Softmax samples are used in backward pass.
Overall, the learnable parameters in the process are $\phi$,$\Lambda$, $\Gamma$, and the estimated adjacency matrix reflecting temporal causal relations is $\sigma(\Lambda)$ and the estimated potential interventional family is $\sigma(\Gamma)$.
@@ 3.2. LLM-guided Meta Initialization
Using tuples provided by LLM, we construct the adjacency matrix $M_{0}^{\mathcal{G}}$ to denote the learned causal structure. These results represent potential causality based on domain information, not guaranteed causal definitions. The subsequent score-based causal discovery optimization ensures the final results conform to causal definitions. We incorporate meta-initialization information as follows:
@@
Guide temporal causal discovery from data
Extract weight
@@
2:  Obtain causal results from LLM and transfer into matrix $M_{0}^{\mathcal{G}}$
@@
7:     Get sub solution $\phi_{t}^{*},\Lambda_{t}^{*},\Gamma_{t}^{*}$, and initialize $\phi_{t+1},\Lambda_{t+1},\Gamma_{t+1}$ by them
@@
11:  Form causal graph by adding an edge whenever $\sigma(\Lambda)>0.5$ is acyclic
@@ 4.1.1. Baselines
  • NeuralGC (Tank et al., 2021): Since NeuralGC only learns contemporaneous relationships, we use $W_{\mathrm{full}}$ defined as below to compress both contemporaneous and time-lagged causal relationships learnt from RealTCD into a single metric for comparison with NeuralGC, which does not differentiate these 2 types.
@@
The distinction in method selection across different settings was intentional and aligned with the capabilities of each algorithm: Since PCMCI and DYNOTEARS are designed for causal structure learning from observational data and can not incorporate interventional data, we included them in the ”Unknown” interventional targets setting. As for TECDI, we use the same data with RealTCD. For baselines that are unable to bring in interventional data, we ensure a fair comparison by keeping the overall sample size consistent across all models, while using only observational data (or normal data in real datasets) for those baselines.
@@ 4.1.2. Synthetic datasets
Before training, all data are normalized by subtracting the mean and dividing by the standard deviation. We experimented on two simulated datasets: Dataset 1 contains 5 nodes, their 1 time-lagged variables and 5 different interventional targets. Dataset 2 contains 10 nodes, their 1 time-lagged variables and 10 different interventional targets. We initially chose to limit our datasets to a time delay of 1 to facilitate the evaluation and presentation. However, it is indeed feasible to increase the time delay of the data.
4.1.3. Real-world data center application
@@
In contemporary data centers, IT equipment stability is crucial. Sophisticated air conditioning systems manage heat, maintaining a consistent temperature. Figure 2 shows a typical data center room organized with rows of equipment (A, B, C, D, etc.), separated by barriers to prevent hot and cold air mixing. Computer Room Air Conditioners (CRACs) on both sides of the room create a closed-loop system to maintain stable conditions. Multiple sensors in the cold aisle provide real-time temperature data, essential for ensuring continuous cold air delivery and stable IT operation.
When an anomaly occurs at a monitoring point, we can find the root cause of the anomaly based on the causal relationships among these entities we have learned through our RealTCD framework.
@@
The data used in this study was obtained from a specific data center at Alibaba. It covers monitoring data from a cooling system of a particular room from January 1st, 2023 to May 1st, 2023, and includes 38 variables in total. These variables comprise 18 cold aisle temperatures from sensors and 20 air conditioning supply temperatures from CRACs. We collected several time series during normal as well as abnormal states. For the latter, data was sampled within 20 minutes of the occurrence of the abnormality, with each sampling interval being 10 seconds. Anomaly points were identified by learning the normal distribution range from historical data, using the $n$-$\sigma$ method. Any data points that fall outside of the $n$-$\sigma$ range (e.g., 3 to 5) of itself are extracted as anomaly time points.
@@ 4.1.4. Evaluation metrics
We leverage two metrics to evaluate the performance of learning causal graph: i) structural Hamming distance (SHD), which calculates the number of different edges (either reversed, missing or redundant) between two DAGs; ii) structural interventional distance (SID), which represents the difference between two DAGs according to their causal inference conditions(Peters and Bühlmann, 2015).
@@
Given the absence of ground truth DAGs in real case, we employ four performance metrics based on expert knowledge to assess algorithms’ efficacy. These metrics mainly consider the physical location relationships within the room, shown in figure 2. i) A2C True counts the correctly identified causal edges from air conditioning supply temperatures (A) to the temperatures of adjacent cold aisles (C), assuming that A influences C. ii) A2A True counts the correctly identified causal edges between adjacent A units, assuming mutual influences among them. iii) C2C True counts the correctly identified causal edges between temperatures in the same column of cold aisles, assuming direct interactions among them. iv) C2A False counts the incorrectly identified causal edges from cold aisle temperatures (C) to air conditioning supply temperatures (A), as such causality is implausible given that downstream variables (C) cannot influence upstream variables (A).
@@ 4.2.1. On synthetic datasets
The results on synthetic datasets are reported in Table 2. On both datasets, our method outperforms baseline models on SHD and SID metrics, with small standard deviations. For the dataset with 10 nodes, the improvement is more pronounced, highlighting our method’s advantages in handling a large number of variables, making optimization more focused and effective. RealTCD consistently achieves better performance in each setting. ”Known” targets setting performs better than the ”Unknown” targets setting since the former employs ground-truth interventional target labels for training. However, since ground-truth targets are often unavailable in real scenarios, RealTCD’s effectiveness with unknown targets is highly practical.
@@
Figure 3(a) and Figure 3(b) display example results of the temporal causal DAG and interventional targets obtained through RealTCD. In Figure 3(b), each row represents a contemporary node, and each column represents a regime corresponding to each interventional family: the first for observational data and the next five for different interventional data. Pink cells indicate the learned or ground truth interventional targets in each regime. These figures demonstrate our method’s ability to accurately identify both the temporal causal graph and interventional targets.
(a) DAG results.
(b) Interventional targets results.
Figure 3. Showcases of the results on synthetic data.